OpenStudy (anonymous):

If m∠7 = 105, what is the measure of ∠10? A. 15° B. 105° C. 75° D. 95

OpenStudy (anonymous):



OpenStudy (anonymous):

∠7 = ∠15 ∠10 = 180 - ∠15 ∠10 = 180 - ∠7 ∠10 = 180 - 105 ∠10 = 75

OpenStudy (anonymous):

If m∠4 = 65, what is the measure of ∠1? A. 25° B. 65° C. 115° D. 135° can you help w. this one too please?

OpenStudy (anonymous):

Same thing, different angles ∠4 = ∠2 ∠1 = 180 - ∠2 ∠1 = 180 - ∠4 ∠1 = 180 - 65 ∠1 = 115

OpenStudy (anonymous):

oh okaay so is that how you set it up always??

OpenStudy (anonymous):

thank you!!